UT 20131114 LUCI2 Nighttime Commissioning

Observer: WSeifert, JHeidt, PBuschkamp, APramskiy, KPolsterer, MEdwards, DThompson
Telescope Operator: SAllanson
LBTO Support: TSargent

Plan for the night

  • AGw transformation verification
  • Center of rotation
  • Test LS-acquisition
  • Stability guiding
  • Preset Bino-Mode, ZPs N1.8/N3.75
  • Test MOS-Acq.
  • Spec-test G200
  • Flats

Summary

Took a long series of images for stability of guiding+large offsets

Took first spectra on sky

Did a successfull bino-preset

AGw transform still not OK, switched to the one obtained with IRTC

LS-acquisition does not work properly

Lost about 2hrs due to telescope issues

Preparations

Michelle again having to work hard to prepare plenty of scripts...

Details

00:30 Skipped flats, since telescope not ready since work going on in dome

00:50 Steve prepares the telescope

01:00 Weather: Clear outside, Temp = +3, Hum = 80% (will hopefully drop at some point)

AGw-transform verification

01:14 Check AGw transform, BS9188, Ks, dither pattern with 10 large 1' offsets , #3-12, stopped

01:30 AGw transform verification as wished by DMiller one StoneO with pointingmask
Stone_O @ PA=0
offset sky 0017
gs1 0018
gs11 failed...this star will not work at PA=0

Stone_O @ PA=-140 sky 0019
gs27 0020
gs30 0021, 22
gs18 0023
gs09 0024
gs22 0025
gs20 0026
gs10 0027
gs12 0028
gs26 0029
gs32 0030
gs24 0031
gs15 0032
gs11 0033

Closed loop ACTIVE on gs26 @ 3 Pas:
0034 PA=180
0035 PA=240
0036 PA=120

TRACK Mode at 4 PAs for open-loop ROTCEN
FC turned on before each image, off after
0037 PA=90
0038 PA=270
0039 PA=180
0040 PA=0

  • ROTCEN: 1031 1048 in TRACK mode

Quick check of the pointing, TRACK mode. SAllanson sent presets
0041 WT10_147 around EL=30
0042 ACT0238 around EL=60
0043 WT10_291 around EL=82
  • Pointing looks OK.

Not sure what images 44-49 are.

0050, 0051 blind filter in (persistence from pointing star visible)

01:42 Walter realigned the field stop

01:46 AGw-transform, #13-17, part of transform check

02:12 GEIRS/ROM hang again when taking image from ROM, Alexander fixed

02:16 Continuing taking AGw transform data, files #xx-33

02:xx Continue check AGw transform, StoneO, PA=180, GS 26 + various PAs

02:50 Center of rotation, #37-40

02:52 Pointing checks, #41-51

Tests LS acquisition

03:22 LS-acq with mask for active flexure compensation, GEIRS-error, fixed and cont'd
Didn't work, offset was not sent to OB execution panel, Alexander/TSargent need to fix it

04:00 BS9107, ZPs JHKKs in N3.75/1.8 (for cross-reference), MOS error, fixed by Peter<>

04:43 LS-acq again, offset-problem fixed, but this time telescope pointing is bad, correct pointing, problem with GS

04:59 Same field with different PA=231, GEIRS error, restart, still problems with GS,
AGw transform apparently poor

05:00 Its completely clear now.

05:40 DMiller/DThompson detected (probably) another sign-error in the AGW, checked that, nope

06:00 Switched back to old IRTC-transform

06:06 Retry LS-acq, PA=0, BS9188, #64 backgr, #65 target, #66 slitview, #67 slitview, #68 ima,
doesn't work, star pos: 967,1001, LS: 1024,1005 => Offset req. 56,3 => Star at 974,1080
=> Moved much in y and almost nothing in x, the opposite was expected
Applied DETXY offset of +30" in x via TCS => moved to the opposite => something wrong

06:47 GEIRS-error, reboot entire LUCI system, #72 starting point, #73 offset +30" RA OK,
#74 offset 0,+30" OK => RADEC offset are correct, DETXY offsets in y

07:12 BS9188_plus, 5 pos, files #75-79 => x-flipped

07:20 => Flipping sign for sending offsets in x within LUCI-SW

07:32 New SW installed, BS9188_plus for verification (#81-85), no flip within x! (misunderstanding by Alexander)
verified via fake LS-acq.

08:00 Revised SW installed, BS9188_plus for verification (#86-90), looks OK

08:10 Did a fake LS-acq => offset are wrongly calculated in alpha (folded by cos (delta))
needs to be checked by Alexander

W3 - test for stability of guiding, PR-type object

09:20 Tests for stability of guiding (pot. PR-target) + large offsets, W3, script W3main_JHK, files #92-104
images in K, H, J, 34 each, among them with a set of sky-images offset by a couple of arcmin
Variable condensation on LUCi2 entrance window, clearly correlated with the ambient hum, nada PR,
seeing degrades from 0.6-1.5" during the observations
WOW, super long script with 104 images went through without problems

Preset binocular mode

10:56 Upload new SW-version, initialize subsystems, authorize LUCI on both sides
11:10 Preset binocular mode, BS9107, failed, TCS claimed that we wanna go below horizon
This is definitively not the cause
11:21 Tried it with a script from Alexander, doesn't work properly
11:30 Did a couple of more tests. In the end the bino-preset was successfull (all subsystems worked),
but the telescope couldn't collimate on left side.

12:27 Preset BS9126, retry LS-acq, change in SW, offsets are now hopefully correctly delivered
Does not work. Acq. ima is #197: Star has x,y=1125.3,1135.8, requested pos: 1278.5,1273.8
calculated offsets: x,y=153,138 (correct), offsets sent to telescope: x,y=-34.56,16.56".
Offset in y is correct, offset in x is wrong. (Sign is OK). With a pixel scale of 0.12"/pix,
the offset should have been -18.36". No idea where it comes from.
We applied offset, resulting file is #198. Position of star is now 1411.8, 1281.25.

First spectra with LUCI2 - Comet Lovejoy

13:00 Comet Lovejoy: collimation nearby, than track preset. Well centered on slit
Took a couple of spectra with G200, lambda_cen=1.93 (3*6sec, 8*60sec),
afterwards spectra of Telluric (check which one)+a couple of flats with lamp2.

13:36 K-band twilight flats, data taken via GEIRS, luci2.20131110.0014-49

13:40 Close dome

-- JochenHeidt - 13 Feb 2014
Topic revision: r11 - 13 Feb 2014, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback